HUBBLE SPACE TELESCOPE DAILY REPORT #1106 PERIOD COVERED: 0700 EST Daily Status Report as of 089/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5565 (WF/PC-2 Cycle 4: Photometric Calibration Monitor: Fields) The WF/PC-2 was used to make photometric calibration observations by taking images of the large globular cluster Omega Centauri (aka, NGC-5139) to help determine astrometric and optical stability of the WF/PC-2. The OSS monitored a total of ten observations, and assessed the data quality. No anomalies were noted, except for recenterings. 1.2 Completed Two Sets of WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxies 3C303.1 and 3C14.1 as part of a continuum survey and to study the galaxies' morphology. Two images were received and displayed. One was okay but encountered some recentering, and the other was executed in fine lock with no apparent problems. 1.3 Completed Two Sets of WF/PC-2 4709 (WF/PC-2 Internal Monitor) The WF/PC-2 camera was used to make a series of routine ultraviolet flat calibration images to monitor the instrument's performance. The observations were monitored in the OSS, and the data quality was assessed. All three observations were reported to be underexposed. 1.4 Completed WF/PC-2 5415 (The Distance to M-96 Galaxy Group: Cycle 4 Medium) The WF/PC-2 was used to make observations of Cepheid variable stars within the galaxy UGC-5889 which itself is located in the M-96 galaxy cluster. This is to help determine the distance to this galactic group, and to improve the determination of the Hubble constant. Four observations were received and displayed. There were FGS recentering during each exposure. In addition the guide star re-acquisition for the last observations was completed late, so the Take Data Flag (TDF) didn't come up until nine seconds into the exposure. The shutter didn't open until the TDF was up, so the exposure was nine seconds short. 1.5 Completed WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earth Flats, Large Set) The WF/PC-2 was used to make a set of Earth flat calibration images using different filters in order to construct high quality flat fields of the OTA illumination pattern. A total of 15 images were received and displayed. One was saturated, five were underexposed, and rest were okay. One observation still remains on the science tape recorder, and will be analyzed later. 1.6 Completed FOC/BL/RD 5626 (SMOV FOS Plate Scale (Revisited - Replaces 5617)) The Faint Object Camera (Both Red and Blue Sides) was used to make plate scale calibration observation of cluster NGC-188-255. The onboard target acquisition was received in real time and left the target well centered. The IMAGE-mode observations were executed with no apparent problems. Complete analysis is pending the science tape recorder dump. No anomalies were noted. 1.7 Completed WF/PC-2 5369 (HST Medium-Deep Survey: Cycle 4 Medium, High Galactic Latitude) The WF/PC-2 made a series of observations of high latitude objects as part of a medium deep sky survey. Several different filters were used. The observations were taken in parallel with other prime science observations. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled :7 Successful :7 2.3 Significant HSTARs: None 2.4 Operations Notes: The FOS/FGS fine alignment (blue) of small apertures was the only SMOV activity accomplished during week #15. SMOV functions still to be accomplished are: FOS/FGS fine alignment (red) of small apertures, FOS plate scale, the GHRS ERO with a supporting WF/PC-2 observation, and a repeat of the FOS ERO to correct a target acquisition/pointing problem. With the exception of the last item which will be scheduled as soon as possible within STScI scheduling guidelines, all SMOV activities are scheduled to be completed by April 3, 1994. For this 24-hour period, each HST subsystem reported trouble-free operations. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of WF/PC-2 5624 (Cycle 4 DD HST Observations of Comet Shoemaker-Levy (1993E) for March) 3.2 Complete Two Sets of WF/PC-2 5570 (WF/PC-2 Cycle 4 Calibration: Earthflats, Large Set) 3.3 Complete WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) 3.4 Complete WF/PC-2 5324 (Globular Cluster Core Imaging) 3.5 Complete WF/PC-2 5415 (The Distance to the M-96 Galaxy Group: Cycle 4 Medium) 3.6 Complete WF/PC-2 5109 (A Deep Survey at High Galactic Latitude (WC06): Cycle 4) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-